OpenStudy (anonymous):

Choose the equation below that represents the line passing through the point (2, -5) with a slope of -3. y = -3x - 13 y = -3x + 11 y = -3x + 13 y = -3x + 1

jimthompson5910 (jim_thompson5910):

they all have a slope of -3, so let's focus on the point (2,-5)

jimthompson5910 (jim_thompson5910):

the line must go through (2,-5), so if you plug it into the correct equation, you'll get a true result

jimthompson5910 (jim_thompson5910):

for instance, in choice A y = -3x - 13 -5 = -3(2) - 13 ... plug in (2, -5) -5 = -6 - 13 -5 = -19 which is FALSE, so (2,-5) is NOT on the line y = -3x - 13

jimthompson5910 (jim_thompson5910):

so that rules out choice A

jimthompson5910 (jim_thompson5910):

see how that all works?